Explain why the average variable cost (AVC) and average total cost (ATC) curves are U-shaped (decreasing, reaches minimum and then increasing)?
1
Expert's answer
2020-09-01T11:10:26-0400

Any manufacturer is interested in how much it costs to produce a unit of output on average. Highlighted average total costs (ATC), average variable costs (AVC) and average fixed costs (AFC).

Average Fixed Cost (AFC) Provides fixed costs per unit of output. They are determined by dividing fixed costs by the amount of output:

"AFC=\frac {FC}{Q}"

As the volume of output increases, the average fixed costs will decrease.

Average Variable Cost (AVC) represents the variable cost per unit of output, and is obtained by dividing the variable cost by the volume of output:


"AVC=\frac {VC}{Q}"

Average Total Cost (ATC)* shows the total cost per unit of production and is determined by the formula:


"ATC=\frac {TC}{Q}"

Since total costs can be represented as the sum of fixed and variable costs (TC = FC + VC), the value of average total costs is determined as the sum of average fixed and average variable costs


"ATC=AFC+AVC"

There are important relationships between marginal, average total and average variable costs. First of all, this concerns the relationship between MC and AVC. If the variable costs per unit of output are higher than the marginal costs, then they decrease with each subsequent unit of output. In the event that the AVC becomes smaller than the MS, then the AVC value starts to increase.